By using mathematical induction prove that (n+1)! > 2^(n+1) for n, where n is a positive integer greater than or equal to 4


1
Expert's answer
2021-02-04T06:52:19-0500

Prove "(n+1)! > 2^{(n+1)}" :


Base case: "n = 4"

"(4+1)! = 5! = 120 > 2^{(4+1)} = 2^5 = 32" -> true


suppose for some n: "n! > 2^n" , let's prove that "(n+1)! > 2^{(n+1)}"

"(n+1)! = n! * (n+1)"

"2^{(n+1)} = 2 * 2^n"

using assumption that "n! > 2^n" it is obvious that "n! * (n+1) > 2 * 2^n" , since "n+1 > 2" for "n \ge 4" .